OpenStudy (anonymous):

hi can you please help me Find the solution of three times the square root of x+6=-12

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

you mean\[3 \times \sqrt{x+6=-12}\]

OpenStudy (anonymous):

or \[\sqrt[3]{x+6=-12}\]